For the point-slope equation given, state the slope and a point on the graph of y+4=-2(x-2) by looking at the equation. From the equation the slope is?
Answer by FrankM(1040)

y+4=-2(x-2)
is of the form point-slope
(y-y1)=m(x-x1)
The slope is -2 and the point is (2,-4)